OpenStudy (anonymous):

given the quadratic function y=x2+2x−3 find the coordinates of the y-ntercept.

OpenStudy (callisto):

Put x =0 into the equation, you'll get the value of y-int

OpenStudy (anonymous):

on y-axis, x = 0 so y int= -3

OpenStudy (callisto):

y=(0)^2+2(0)−3 = 0+0 -3 =?

OpenStudy (anonymous):

y=-3

OpenStudy (anonymous):

and for the x-intercept would you put it in as -3=x^2+2x-3

OpenStudy (callisto):

y-int -> yes For x-int, put y=0 and solve the equation

OpenStudy (anonymous):

ok im kinda confused

OpenStudy (callisto):